\providecommand{\noopsort}[1]{}
\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
% \MRhref is called by the amsart/book/proc definition of \MR.
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}
\begin{thebibliography}{AFMO24b}

\bibitem[AFMO24a]{HAI}
Jeffrey~D. Adler, Jessica Fintzen, Manish Mishra, and Kazuma Ohara,
  \emph{\noopsort{1}{S}tructure of {H}ecke algebras arising from types}, 2024,
  preprint available at https://arxiv.org/abs/2408.07805.

\bibitem[AFMO24b]{HAIKY}
\bysame, \emph{\noopsort{2}{R}eduction to depth zero for tame $p$-adic groups
  via {H}ecke algebra isomorphisms}, 2024, preprint available at
  https://arxiv.org/abs/2408.07801.

\bibitem[BK98]{BK-types}
Colin~J. Bushnell and Philip~C. Kutzko, \emph{Smooth representations of
  reductive {$p$}-adic groups: structure theory via types}, Proc. London Math.
  Soc. (3) \textbf{77} (1998), no.~3, 582--634. \MR{1643417}

\bibitem[Fin]{Fintzen-IHES}
Jessica Fintzen, \emph{Supercuspidal representations: construction,
  classification, and characters}, preprint available at
  \url{https://www.math.uni-bonn.de/people/fintzen/IHES_Fintzen.pdf}.

\bibitem[Fin21]{Fi-exhaustion}
\bysame, \emph{Types for tame {$p$}-adic groups}, Ann. of Math. (2)
  \textbf{193} (2021), no.~1, 303--346. \MR{4199732}

\bibitem[FKS23]{FKS}
Jessica Fintzen, Tasho Kaletha, and Loren Spice, \emph{A twisted {Y}u
  construction, {H}arish-{C}handra characters, and endoscopy}, Duke Math. J.
  \textbf{172} (2023), no.~12, 2241--2301. \MR{4654051}

\bibitem[GR05]{GoldbergRoche-Hecke}
David Goldberg and Alan Roche, \emph{Hecke algebras and {${\rm SL}_n$}-types},
  Proc. London Math. Soc. (3) \textbf{90} (2005), no.~1, 87--131. \MR{2107039}

\bibitem[KY17]{Kim-Yu}
Ju-Lee Kim and Jiu-Kang Yu, \emph{Construction of tame types}, Representation
  theory, number theory, and invariant theory, Progr. Math., vol. 323,
  Birkh\"{a}user/Springer, Cham, 2017, pp.~337--357. \MR{3753917}

\bibitem[{May}20]{2020arXiv200106259M}
Arnaud {Mayeux}, \emph{{Comparison of Bushnell-Kutzko and Yu's constructions of
  supercuspidal representations}}, arXiv e-prints (2020), arXiv:2001.06259v2.

\bibitem[Mor93]{Morris}
Lawrence Morris, \emph{Tamely ramified intertwining algebras}, Invent. Math.
  \textbf{114} (1993), no.~1, 1--54. \MR{1235019}

\bibitem[MP94]{MR1253198}
Allen Moy and Gopal Prasad, \emph{Unrefined minimal {$K$}-types for {$p$}-adic
  groups}, Invent. Math. \textbf{116} (1994), no.~1-3, 393--408. \MR{1253198}

\bibitem[MP96]{MR1371680}
\bysame, \emph{Jacquet functors and unrefined minimal {$K$}-types}, Comment.
  Math. Helv. \textbf{71} (1996), no.~1, 98--121. \MR{1371680}

\bibitem[Roc02]{Roche:parabolic}
Alan Roche, \emph{Parabolic induction and the {B}ernstein decomposition},
  Compositio Math. \textbf{134} (2002), no.~2, 113--133. \MR{1934305}

\bibitem[Yu01]{Yu}
Jiu-Kang Yu, \emph{Construction of tame supercuspidal representations}, J.
  Amer. Math. Soc. \textbf{14} (2001), no.~3, 579--622. \MR{1824988}

\end{thebibliography}

\documentclass{amsart}
\usepackage{amsthm}
\usepackage{enumerate}
\usepackage{amsmath,amssymb}
\usepackage{tikz-cd}

\usepackage{xr-hyper} 

\usepackage{hyperref}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}
\theoremstyle{definition}
\newtheorem{remark}[theorem]{Remark}
\numberwithin{equation}{section}
\allowdisplaybreaks[1]

\DeclareMathOperator{\End}{End}
\DeclareMathOperator{\Hom}{Hom}

\DeclareMathOperator{\Res}{Res}
\DeclareMathOperator{\ord}{ord}
\DeclareMathOperator{\aff}{aff}

\DeclareMathOperator{\GL}{GL}
\DeclareMathOperator{\SL}{SL}
\DeclareMathOperator{\prim}{prim}
\DeclareMathOperator{\Rep}{Rep}
\newcommand{\IEC}{\mathfrak{I}}
\DeclareMathOperator{\sgn}{sgn}
\DeclareMathOperator{\Lie}{Lie}
\DeclareMathOperator{\Gal}{Gal}
\DeclareMathOperator{\norm}{norm}
\DeclareMathOperator{\Tr}{Tr}

\renewcommand{\restriction}{|}

\newcommand{\tG}{{\widetilde G }}
\newcommand{\tM}{{\widetilde M }}

\newcommand{\tT}{{\widetilde T }}
\newcommand{\Krel}{{\mathcal{K}\textup{-rel}}}
\newcommand{\flength}{{\ell_{\Krel}}}
\newcommand{\isoarrow}{\stackrel{\sim}{\longrightarrow}}
\newcommand{\muT}{\mu^{\cT}}

\newcommand{\Wzero}{\Omega(\rho_{M})}
\newcommand{\Wzeroz}{\Omega(\rho_{M^0})}
\newcommand{\plength}{\ell_{\prim}}
\newcommand{\phm}{\phantom{-}}

\newcommand{\bC}{\mathbb{C}}
\newcommand{\bZ}{\mathbb{Z}}

\newcommand{\cB}{\mathcal{B}}
\newcommand{\cH}{\mathcal{H}}
\newcommand{\cO}{\mathcal{O}}
\newcommand{\cT}{\mathcal{T}}

\newcommand{\ff}{\mathfrak{f}}
\newcommand{\fp}{\mathfrak{p}}
\newcommand{\fs}{{\mathfrak{s}}}

\title[A depth-zero principal-series block...]%
{A depth-zero principal-series block whose Hecke algebra has a non-trivial two-cocycle}
\author{Jeffrey D. Adler}
\address{Department of Mathematics and Statistics\\
        American University\\
        4400 Massachusetts Ave NW\\
        Washington, DC 20016-8050, USA}
\email{jadler@american.edu} 
\author{Jessica Fintzen}
\address{Universität Bonn\\
	Mathematisches Institut\\
	Endenicher Allee 60\\
	53115 Bonn\\
	Germany}
\email{fintzen@math.uni-bonn.de}
\author{Kazuma Ohara}
\address{Max Planck Institute for Mathematics\\ Vivatsgasse 7, 53115 Bonn, Germany}
\email{kazuma@mpim-bonn.mpg.de}

\begin{document}

\externaldocument[HAI-][nocite]{Adler--Fintzen--Mishra--Ohara_Structure_of_Hecke_algebras_arising_from_types}[https://arxiv.org/pdf/2408.07801]
\externaldocument[HAIKY-][nocite]{Adler--Fintzen--Mishra--Ohara_Reduction_to_depth_zero_for_tame_p-adic_groups_via_Hecke_algebra_isomorphisms}[https://arxiv.org/pdf/2408.07805]

\begin{abstract}
Recently the authors have shown that every Hecke algebra
associated to a type constructed by Kim and Yu is isomorphic to a Hecke algebra
for a depth-zero type. 
An example in the literature has been suggested
as a counterexample to this result.
We show that the example is not a counterexample, and exhibit
some of its interesting properties, e.g., we show that a principal series, depth-zero type can have a Hecke algebra with non-trivial two-cocyle, a phenomenon that many did not expect could occur.
\end{abstract}

\maketitle

\section{Introduction}
Let $G$ denote a connected reductive group over a non-archimedean local field $F$.
The category $\Rep(G(F))$ of smooth, complex representations of $G(F)$
is a direct product of full subcategories
called ``Bernstein blocks'':
\[
\Rep(G(F))=\prod_{\fs\in\IEC(G)}\Rep^\fs(G(F)).
\]
Each of the blocks $\Rep^\fs(G(F))$
is equivalent to the category of unital right modules
over an algebra $\cH^\fs$.
Suppose that the category $\Rep^\fs(G(F))$ has
an associated \emph{type}, as defined by Bushnell and Kutzko
\cite{BK-types}, i.e., a compact open subgroup $K$ of $G(F)$
and an irreducible smooth representation $\rho$ of $K$ such that
a representation $\pi \in \Rep(G(F))$ belongs to $\Rep^\fs(G(F))$
if and only if every irreducible subquotient of $\pi$
contains $\rho$ upon restriction to $K$.
Then we can replace the algebra $\cH^\fs$
by the Hecke algebra $\cH(G(F),(K, \rho))$
of all compactly supported,
$\End_{\bC}(\rho)$-valued
functions on $G(F)$
that transform on the left and right
according to $\rho$.
That is, $\Rep^\fs(G(F))$ is equivalent to the category
of modules over $\cH(G(F),(K, \rho))$.

One of the present authors \cite{Fi-exhaustion} has shown that,
provided that $G$ splits over a tamely ramified extension of $F$ and the residual characteristic $p$ of $F$ is not too small,
the construction of Kim and Yu \cite{Kim-Yu}
provides types for every Bernstein block for $G(F)$.
Thus, under this mild tameness assumption, one can in principle understand
the category $\Rep(G(F))$ by understanding the structures
of all of the Hecke algebras that arise from the types constructed by Kim and Yu.

In the ``depth-zero'' case, the compact group $K$
contains a parahoric subgroup of $G(F)$, and the representation
$\rho$ is trivial on the pro-$p$ radical of this parahoric.
In the special case where $K$ is a parahoric subgroup,
Morris \cite[Theorem 7.12]{Morris} has described the structures of these Hecke algebras.
More generally, the authors
\cite[Theorem \ref{HAI-theoremstructureofheckefordepthzero}]{HAI}
have described the structures of all depth-zero types.

In the set-up of Kim and Yu (\cite{Kim-Yu}), a type $(K,\rho)$ for $G$ of positive depth
is constructed from a depth-zero type $(K^0,\rho^0)$ for a subgroup $G^0$ of $G$,
together with some additional data.
The authors have recently shown
\cite[Theorem \ref{HAIKY-theoremstructureofheckeforKimYu}]{HAIKY}
that the associated Hecke algebras
$\cH(G(F), (K, \rho))$ and $\cH(G^0(F),(K^0, \rho^0))$ are isomorphic, after twisting the construction by Kim and Yu by a quadratic character arising from \cite{FKS},
thus describing 
the structures of all Hecke algebras arising from Kim--Yu types.

In outline,
from the pair $(K,\rho)$, one constructs
a group $W^\heartsuit$,
and
a normal, affine reflection subgroup $W_{\aff}$ of $W^\heartsuit$.
Choosing a set $S$ of generating reflections for $W_{\aff}$,
one constructs a parameter function $q\colon S \longrightarrow \bC^\times$,
thus obtaining an abstract Hecke algebra $\cH(W_{\aff},q)$.
The choice of $S$ gives rise to a complement $\Omega$
to $W_{\aff}$ in $W^\heartsuit$.
A choice of a family $\cT$ of intertwining operators
gives rise to a $2$-cocycle
$\muT \colon \Omega \times \Omega \longrightarrow \bC^\times$.
One then obtains an isomorphism of $\bC$-algebras
\[
\cH(G(F),(K, \rho)) \isoarrow \bC[\Omega,\muT] \ltimes \cH(W_{\aff},q).
\]
That is, $\cH(G(F),(K, \rho))$ is isomorphic to a semidirect product
of our abstract Hecke algebra and the $\muT$-twisted group algebra of $\Omega$,
where the structure of multiplication between these two factors is controlled
by the conjugation action of $\Omega$ on $W_{\aff}$.

In relation to the above discussion, Roche \cite[\S4]{Roche:parabolic} and
Goldberg--Roche \cite[\S11.8]{GoldbergRoche-Hecke}
each illustrate some unusual phenomena by presenting
an example, that they attribute to Kutzko,
of a Hecke algebra $\cH := \cH(G(F), (K, \rho))$ for a particular block of $G = \SL_{8}$.
In this note,
we discuss this example, determine an attached depth-zero pair $(K^0, \rho^0)$ and describe the depth-zero algebra
$\cH^0 := \cH(G^0(F),(K^0, \rho^0))$
that corresponds to it via 
\cite[Theorem \ref{HAIKY-theoremstructureofheckeforKimYu}]{HAIKY} explicitly, as well as the closely related Hecke  algebra $\cH^{0,\circ} :=\cH(G^0(F),(K^{0, \circ}, \rho^0))$, where we replace $K^0$ by the parahoric subgroup $K^{0,\circ}$ contained in it.
We have several aims in doing so.
\begin{enumerate}[(a)]
\item
First, 
a
remark in \cite{GoldbergRoche-Hecke} that
$\cH$ cannot be isomorphic
to any of the intertwining algebras constructed by
Morris \cite{Morris} let several mathematicians believe that $\cH$ would provide a counterexample to 
\cite[Theorem \ref{HAIKY-theoremstructureofheckeforKimYu}]{HAIKY}.
Thus, we want to assure readers that this is not the case.
\item
Second, $\cH^0$ 
provides an example of a depth-zero Hecke algebra
where the associated affine reflection group is trivial,
the group $\Wzero$ is nonabelian and infinite,
and the cocycle $\muT$ is non-trivial.
In particular, $\cH^0$ is an example of a Hecke algebra attached to a depth-zero, principal-series block of a quasi-split group that requires a non-trivial 2-cocycle, something that was long believed not to exist. 
We believe that this example might be useful
for researchers in the future.
\end{enumerate}

\subsection*{Notation}

For a connected reductive group $G$ and a reductive subgroup $M$ of $G$, let $N_{G}(M)$, resp., $Z_{G}(M)$, denote the normalizer, resp., centralizer, of $M$ in $G$.

For a finite field extension $E/F$ and $A$ a linear algebraic group or a Lie algebra thereof defined over $E$, we write $\Res_{E/F}(A)$ for the Weil restriction of $A$ to $F$.

For a linear algebraic group $G$, we denote by $\Lie(G)$ the Lie algebra of $G$ and by $\Lie^{*}(G)$ the dual of $\Lie(G)$.
We also write $\Lie^{*}(G)^{G}$ for the subscheme of $\Lie^{*}(G)$ fixed by the coadjoint action of $G$ on $\Lie^{*}(G)$.
For a morphism $f \colon G \to H$ of algebraic groups, let 
\[
\Lie(f) \colon \Lie(G) \to \Lie(H)
\]
denote the morphism between their Lie algebras induced by $f$.

Suppose that $G$ is a connected reductive group defined over a non-archimedean local field $F$.
We denote by $\mathcal{B}(G, F)$ the enlarged Bruhat–Tits building of $G$.
For $x \in \mathcal{B}(G, F)$, let $G(F)_{x}$ denote the stabilizer of $x$ in $G(F)$.
For $r \in \mathbb{R}$ with $r \ge 0$, we also let $G(F)_{x, r}$, resp., $G(F)_{x, r+}$, be the Moy--Prasad filtration subgroup of $G(F)$ of depth $r$, resp., $r+$, associated to $x$ (see \cite{MR1253198, MR1371680}).
We use the analogous notation for the Lie algebra $\Lie(G)$ and its dual $\Lie^{*}(G)$, where $r$ is allowed to be any element of $\mathbb{R}$.

For a compact, open subgroup $K$ of $G(F)$ and an irreducible smooth representation $\rho$ of $K$, we denote by $\cH(G(F), (K, \rho))$ the Hecke algebra attached to $(K, \rho)$.
We refer to \cite[Section \ref{HAI-Hecke algebras and endomorphism algebras}]{HAI} for the precise definition of $\cH(G(F), (K, \rho))$.

Suppose that $K$ is a subgroup of a group $H$ and $h \in H$. We denote $hKh^{-1}$ by $^hK$. If $\rho$ is a representation of $K$, we write $^h\!\rho$ for the representation $x\mapsto \rho(h^{-1}xh)$ of $^hK$.

\section{The example}
In this section we introduce the example studied in this paper that we learned about from
Roche \cite[\S4]{Roche:parabolic} and
Goldberg--Roche \cite[\S11.8]{GoldbergRoche-Hecke},
who attribute it to Kutzko.
We present it, a type for the group $\SL_8$, in the language of Kim and Yu's construction of types.
Doing so then allows us to describe the associated depth-zero type for a smaller group $G^0$,
seeing directly that it fits into our set-up.

Let $F$ denote a non-archimedean local field with residue field $\ff$
of characteristic $p$ (assumed odd) and order $q$.
We fix a uniformizer $\varpi_F$ of $F$
and a square root $\sqrt{-1}$ of $-1$ in $\bC^\times$.
For any finite field extension $E/F$,
we denote by $\cO_E$ the ring of integers in $E$,
by $\fp_E$ the prime ideal in $\cO_E$, by $\Tr_{E/F} \colon E \to F$ the trace map, and by
$N_{E/F} \colon E^\times \to F^\times$
 the norm map.
Let $\ord$ denote the discrete valuation on $F^{\times}$ with the value group $\mathbb{Z}$.
 For any finite extension $E$ of $F$ , we also write $\ord$ for the unique extension of this valuation to $E^{\times}$.
We denote by $\ord^{\norm}_E \colon E^{\times} \twoheadrightarrow \bZ$ the normalized valuation on $E^{\times}$.

Let $\zeta$ be a primitive $(q - 1)$-st root of unity in $F$.
Assume that $4$ divides $q - 1$.
It follows that there exists a unique character $\eta \colon F^{\times} \to \mathbb{C}^{\times}$ that is trivial on $\varpi_{F}$ and $1 + \fp_{F}$ and satisfies $\eta(\zeta) = \sqrt{-1}$.
Let $E_{2}$ be the splitting field of the polynomial $X^{2} + \varpi_{F}$,
and $E_{4}$ the splitting field of the polynomial $X^{4} + \zeta \varpi_{F}$.
(Note that the fields that we denote by $E_{2}$ and $E_{4}$ here are denoted by $E_{1}$ and $E_{2}$, respectively, in \cite{Roche:parabolic}.)
Let $\varpi_{E_{2}}$, resp., $\varpi_{E_{4}}$,
denote a uniformizer of $E_{2}$, resp., $E_{4}$,
such that $\varpi_{E_{2}}^{2} = - \varpi_{F}$, resp., $\varpi_{E_{4}}^{4} = - \zeta \varpi_{F}$.
We fix a generator $\sigma_{2}$ of the Galois group $\Gal(E_2/F)$ and a generator $\sigma_{4}$ of the Galois group $\Gal(E_{4}/F)$.

We define the following reductive groups over $F$:
\begin{align*}
\tG  = \tG^2 &= \GL_8, \\
\tG^1 &= \Res_{E_2/F} (\GL_2) \times \GL_4, \\
\tG^0 &= \Res_{E_2/F} (\GL_2) \times \Res_{E_4/F} (\GL_1), \\
\tT = \tM^0 
&= \Res_{E_2/F} \left(
\GL_1 \times \GL_1
\right) \times \Res_{E_4/F} (\GL_1).
\end{align*}
We identify $\GL_{1} \times \GL_{1}$ with the diagonal torus of $\GL_{2}$ by the map
\(
(t_{1}, t_{2}) \mapsto 
(\begin{smallmatrix}
t_{1} & 0 \\
0 & t_{2}
\end{smallmatrix}),
\)
thus obtaining an
embedding $\tM^0 \hookrightarrow \tG^0$.
Fix isomorphisms
$F^{\oplus 2} \cong E_2$
and
$F^{\oplus 4} \cong E_4$
of vector spaces over $F$,
thus determining isomorphisms
$$
F^{\oplus 8}
\quad \cong \quad
E_2^{\oplus 2} \oplus F^{\oplus 4}
\quad \cong \quad
E_2^{\oplus 2} \oplus E_4.
$$
These choices determine embeddings of $F$-groups
\(
\tG^0
\hookrightarrow \tG^1
\hookrightarrow \tG^2.
\)
Let us identify each group above with its images under these maps,
so that they are all contained within $\tG = \GL_8$.

The maximal split subtorus $A_{\tT}$ of $\tT$
is isomorphic to $\GL_1 \times \GL_1 \times \GL_1$.
Note that $\tM^0 = Z_{\tG^0}(A_{\tT})$.
For $i=1,2$, let
$\tM^i = Z_{\tG^i}(A_{\tT})$,
and write $\tM = \tM^2$.

Let $G = \SL_8$.
For $X \in \{G^i, M^i, M, T \, | \, i = 0, 1, 2\}$,
we let 
$X = \widetilde X \cap G$.
We thus obtain a twisted Levi sequence
$(G^0 \subset G^1 \subset G^2 = \SL_8)$,
and a Levi subgroup $M^0 \subset G^0$.
We denote by $\Phi(X, T)$ the absolute root system of $X$ with respect to the maximal torus $T$.
For $\alpha \in \Phi(X, T)$, we denote by $\alpha^{\vee}$ the corresponding (absolute) coroot.

We let $\widetilde{K}^{0}$ be the Iwahori subgroup of $\tG^0(F)$ given by $\widetilde{K}^{0} = I_{2} \times I_{4}$, where $I_{2}$ denotes the Iwahori subgroup of $\GL_{2}(E_{2}) = \left(
\Res_{E_{2}/F}(\GL_{2})
\right)(F)$ defined by
\[
I_{2} =
\biggl\{
\begin{pmatrix}
a & b \\
c & d
\end{pmatrix} \in \GL_{2}(E_{2})
\biggr. \,\biggl|\,
a, d \in \cO_{E_{2}}^{\times}, b \in \cO_{E_{2}}, c \in \fp_{E_{2}}
\biggr\},
\]
and $I_{4}$ denotes the Iwahori subgroup of $\left(
\Res_{E_{4}/F}(\GL_{1})
\right)(F) = E_4^{\times}$, i.e., $I_{4} = \cO_{E_{4}}^{\times}$. 

We choose $x_{0} \in \cB(M^{0}, F)$ and fix a commutative diagram $\{ \iota \}$
\[
\begin{tikzcd}
	\cB(G^0, F)
	\arrow[r, hook]
	\arrow[dr, pos=.5, phantom, "\circlearrowleft"]
	&
	\cB(G^1, F)
	\arrow[r, hook]
	\arrow[dr, pos=.5, phantom, "\circlearrowleft"]
	&
	\cB(G^2, F)
	\\
	\cB(M^0, F)
	\arrow[u, hook]
	\arrow[r, hook]
	&
	\cB(M^1, F)
	\arrow[u, hook]
	\arrow[r, hook]
	&
	\cB(M^2, F)
	\arrow[u, hook]
\end{tikzcd} 
\]
of admissible embeddings of buildings that is $(0, \tfrac18, \tfrac14)$-generic relative to $x_{0}$ in the sense of \cite[3.5~Definition]{Kim-Yu} such that $G^0(F)_{x_0}=\widetilde K^0 \cap G^0(F)$.
Here and from now on, we identify a point in $\cB(M^{0}, F)$ with its images via the embeddings $\{ \iota \}$. 
Then we have
\[
G^0(F)_{x_{0}, 0} = \left\{
(g_2, g_4) \in (I_{2} \times I_{4}) \cap G^0(F) \mid (\det(g_2) \bmod \mathfrak{p}_{E_{2}}) \cdot (g_{4} \bmod \mathfrak{p}_{E_{4}})^{2} = 1
\right\},
\]
where we regard $(\det(g_2) \bmod \mathfrak{p}_{E_{2}})$ and $(g_{4} \bmod \mathfrak{p}_{E_{4}})$ as elements of $\mathfrak{f}^{\times}$.
Let $K^0$ be either $G^{0}(F)_{x_0}$ or $G^0(F)_{x_0, 0}$.
We also define $\widetilde{K}_{M^0} = \widetilde{K}^0 \cap \widetilde{M}^0(F)$ and $K_{M^0} = K^{0} \cap M^0(F)$.
Thus, we have $\widetilde{K}_{M^0} = \cO_{E_{2}}^{\times} \times \cO_{E_{2}}^{\times} \times \cO_{E_{4}}^{\times}$ and
$K_{M^0} = M^0(F)_{x_0}$ or $M^0(F)_{x_0,0}$
according as $K^0 = G^0(F)_{x_0}$ or $G^0(F)_{x_0,0}$.
More precisely, 
\begin{equation}
\label{descriptionofKM0}
K_{M^0} = 
\begin{cases}
\left\{
(x, y, z) \in \cO_{E_{2}}^{\times} \times \cO_{E_{2}}^{\times} \times \cO_{E_{4}}^{\times}
\mid
N_{E_{2}/ F}(xy) N_{E_{4}/ F}(z) = 1 
\right\} \\
\hspace{20em}\text{if $K^0 = G^0(F)_{x_0}$,}
\\[1ex]
\left\{
(x, y, z) \in \cO_{E_{2}}^{\times} \times \cO_{E_{2}}^{\times} \times \cO_{E_{4}}^{\times}
\;\middle|\;
\begin{aligned}
& N_{E_{2}/F}(xy)\, N_{E_{4}/F}(z) = 1, \\
& (xy \bmod \mathfrak{p}_{E_{2}})\cdot (z \bmod \mathfrak{p}_{E_4})^{2} = 1
\end{aligned}
\right\}
\\
\hspace{20em}\text{if $K^0 = G^0(F)_{x_0,0}$.}
\end{cases}
\end{equation}
We observe that
\begin{equation}
\label{connectedparahoricvsdisconenctedparahoric}
K^{0} = K_{M^0} \cdot G^0(F)_{x_0, 0}.
\end{equation}
Since the embedding $\iota \colon \cB(M^0, F) \hookrightarrow \cB(G^0, F)$ is $0$-generic relative to $x_{0}$, the inclusion $M^0(F)_{x_0,0} \subset G^0(F)_{x_0,0}$ induces an isomorphism 
\[
M^0(F)_{x_0, 0}/M^0(F)_{x_{0}, 0+} \isoarrow G^0(F)_{x_0, 0}/G^0(F)_{x_0, 0+}.
\]
Combining this with \eqref{connectedparahoricvsdisconenctedparahoric}, we also have
\begin{equation}
\label{isomofreductivequotients}
K_{M^0}/M^0(F)_{x_{0}, 0+} \isoarrow K^0/G^0(F)_{x_0, 0+}.
\end{equation}

We define the character $\widetilde\rho_{M^0}$ of $\widetilde{K}_{M^0}$ by $\widetilde\rho_{M^0} = 1 \boxtimes \left(
\eta \circ N_{E_{2} / F}
\right) \boxtimes 1$, and write $\rho_{M^0} = \widetilde\rho_{M^0} \restriction_{K_{M^0}}$.
We define the character $\rho^0$ of $K^0$ as the composition of the surjection $K^0 \twoheadrightarrow K^0/G^0(F)_{x_{0}, 0+}$, the inverse of the isomorphism in \eqref{isomofreductivequotients} and the character $\rho_{M^0}$.
More precisely, $\rho^0$ is the restriction to $K^0$ of the character $\eta_{2} \boxtimes 1$ of the group $\widetilde K^0$, where $\eta_{2}$ denotes the character of $I_{2}$ defined by
\[
\eta_{2} \left(
\begin{pmatrix}
a & b \\
c & d
\end{pmatrix}
\right) = \left(
\eta \circ N_{E_{2} / F}
\right)(d).
\]

Let $E = E_{2}$ or $E_{4}$.
We fix an additive character $\Psi \colon F \rightarrow \bC^{\times}$ that is trivial on $\fp_{F}$ and non-trivial on $\cO_{F}$.
We define a character $\phi_{E}$ of $1 + \fp_{E}$ by $\phi_{E}(1 + x) = \Psi(\Tr_{E/F}(\varpi_{E}^{-1} x))$ for $x \in \fp_{E}$.
We fix an extension of $\phi_{E}$ to $E^{\times}$ and use the same notation $\phi_{E}$ for it.
We also define the character $\phi_{\GL_2(E_2)}$ of $\GL_{2}(E_{2})$ by $\phi_{\GL_2(E_2)}(g) = \phi_{E_{2}}(\det(g))$.
We define the character $\widetilde{\phi}_{0}$ of $\tG^0(F) = \GL_{2}(E_{2}) \times E_{4}^{\times}$ by $\widetilde{\phi}_{0} = 1 \boxtimes \phi_{E_{4}}$, and define the character $\widetilde{\phi}_{1}$ of $\tG^1(F) = \GL_2(E_2) \times \GL_4(F)$ by $\widetilde{\phi}_{1} = \phi_{\GL_2(E_2)} \boxtimes 1$.
We write $\phi_{0} = \widetilde{\phi}_{0} \restriction_{G^0(F)}$ and $\phi_{1} = \widetilde{\phi}_{1} \restriction_{G^1(F)}$.
\begin{lemma}
The character $\phi_{0}$ is $(G^{1}, G^{0})$-generic of depth $\tfrac14$ relative to the point $x_{0}$, and the character $\phi_{1}$ is $(G^{2}, G^{1})$-generic of depth $\tfrac12$ relative to the point $x_{0}$ in the sense of \cite[Definition~3.5.2]{Fintzen-IHES}.
\end{lemma}
\begin{proof}
By construction, $\phi_0$ is trivial on $G^0(F)_{x_0,(1/4)+}$, and  $\phi_1$ is trivial on $G^1(F)_{x_0,(1/2)+}$.
We define $\widetilde X_{0}^{*} \in \Lie^{*}(\tG^{0})^{\tG^0}(F)$ and $\widetilde X_{1}^{*} \in \Lie^{*}(\tG^{1})^{\tG^1}(F)$ as follows.
Let $E \in \{E_2, E_4\}$.
We use the same notation 
\[
\Tr_{E/F} \colon \Res_{E/F}(\Lie(\GL_1)) \to \Lie(\GL_1)
\]
for the usual trace morphism whose map on $F$-valued points is the trace map.
Let
\[
m(\varpi_{E}^{-1}) \colon \Lie\left(
\Res_{E/F}(\GL_{1})
\right) \to \Lie\left(
\Res_{E/F}(\GL_{1})
\right)
\]
denote the morphism induced by multiplication by $\varpi_{E}^{-1} \in E = \Lie\left(
\Res_{E/F}(\GL_{1})
\right)(F)$.
We define $\widetilde X_{0}^{*} \in \Lie^{*}(\tG^{0})^{\tG^0}(F)$ as the composition of the projection map
\[
\Lie(\tG^{0}) \rightarrow \Lie\left(
\Res_{E_{4}/F}(\GL_{1})
\right)
\]
and
\[
\Tr_{E_{4}/F}
 \circ m(\varpi_{E_{4}}^{-1}) \colon \Lie\left(
\Res_{E_{4}/F}(\GL_{1})
\right) \to \Lie(\GL_{1}).
\]
To define $\widetilde X_{1}^{*}$, we let 
\[
\Res_{E_{2}/F}(\det) \colon \Res_{E_{2}/F}(\GL_{2}) \to \Res_{E_2/F}(\GL_{1})
\]
be the morphism of algebraic groups induced by the usual determinant map $\det \colon \GL_{2} \to \GL_{1}$.
Now, we define $\widetilde X_{1}^{*} \in \Lie^{*}(\tG^{1})^{\tG^1}(F)$ as the composition of the projection map
\[
\Lie(\tG^{1}) \rightarrow \Lie\left(
\Res_{E_{2}/F}(\GL_{2})
\right)
\]
and
\[
\Tr_{E_{2}/F}
 \circ m(\varpi_{E_{2}}^{-1})  \circ \Lie\left(
\Res_{E_{2}/F}(\det)
\right) \colon \Lie\left(
\Res_{E_{2}/F}(\GL_{2})
\right) \to \Lie(\GL_{1}).
\]
We define $X_{0}^{*} \in \Lie^{*}(G^{0})^{G^0}(F)$ and $X_{1}^{*} \in \Lie^{*}(G^{1})^{G^1}(F)$ as the restrictions of $\widetilde X_{0}^{*}$ and $\widetilde X_{1}^{*}$ to $\Lie(G^0)$ and $\Lie(G^1)$, respectively. 
Then the restriction of $\phi_{0}$ to 
\[
G^{0}(F)_{x_{0}, 1/4}/G^0(F)_{x_0,(1/4)+} \simeq \Lie(G^0)(F)_{x_{0}, 1/4}/\Lie(G^0)(F)_{x_{0}, (1/4)+}
\]
 is given by $\Psi \circ X_{0}^{*}$, and the restriction of $\phi_{1}$ to 
\[
G^{1}(F)_{x_{0}, 1/2}/G^1(F)_{x_0,(1/2)+} \simeq \Lie(G^1)(F)_{x_{0}, 1/2}/\Lie(G^1)(F)_{x_{0}, (1/2)+}
\]
 is given by $\Psi \circ X_{1}^{*}$.

We will prove that $X_{0}^{*}$ is $(G^1, G^0)$-generic of depth $- 1/4$, and $X_{1}^{*}$ is $(G^2, G^1)$-generic of depth $-1/2$ in the sense of \cite[Definition~3.5.2]{Fintzen-IHES}.
First, we will prove that $X_{0}^{*}$ satisfies {\bf (GE0)} and {\bf (GE1)} in \cite[Definition~3.5.2]{Fintzen-IHES}.
Let $\alpha \in \Phi(G^1, T) \smallsetminus \Phi(G^{0}, T)$.
Then we have
\[
X^{*}_{0}(\Lie( \alpha^{\vee})(1)) = \sigma^{i}_{4}(\varpi_{E_{4}}^{-1}) - \sigma^{j}_{4}(\varpi_{E_{4}}^{-1})
\]
for some $i, j \in \{0, 1, 2, 3\}$ with $i \neq j$.
Since $E_{4} = F[\varpi_{E_{4}}^{-1}]$, we obtain from \cite[Proposition~5.9]{2020arXiv200106259M} that
\[
\ord\left(
\sigma^{i}_{4}(\varpi_{E_{4}}^{-1}) - \sigma^{j}_{4}(\varpi_{E_{4}}^{-1})
\right) = \ord(\varpi_{E_{4}}^{-1}) = -1/4.
\]
Thus, the element $X_{0}^{*}$ satisfies {\bf (GE1)} in \cite[Definition~3.5.2]{Fintzen-IHES}.
Moreover, since $(0, \varpi_{E_{4}}) \in \Lie(G^0)_{x_{0}, 1/4}$ (where we view $\varpi_{E_{4}}$ in $\Lie(\Res_{E_4/F}(\GL_1))(F)$ by identifying the latter with $E_4$ and note that $(0, \varpi_{E_{4}}) \in \Lie(G^0)(F)$ as $\varpi_{E_4}$ has trace zero) and 
\[
\ord\left(
X^{*}_{0}(0, \varpi_{E_{4}})
\right) = \ord(4) = 0,
\]
we have $X^{*}_{0} \not \in \Lie^*(G^0)_{x_{0}, (-1/4)+}$.
Since it can be checked from the definition that $X_{0}^{*} \in \Lie^*(G^0)_{x_0, -1/4}$, the element $X_{0}^{*}$ also satisfies {\bf (GE0)} in \cite[Definition~3.5.2]{Fintzen-IHES}.

Next, we will prove that $X_{1}^{*}$ satisfies {\bf (GE0)} and {\bf (GE1)} in \cite[Definition~3.5.2]{Fintzen-IHES}.
Let $\alpha \in \Phi(G^2, T) \smallsetminus \Phi(G^{1}, T)$.
Then, using that $\sigma_2(\varpi_{E_2})=-\varpi_{E_2}$, we obtain
\[
X^{*}_{1}(\Lie(\alpha^{\vee})(1)) \in \{\pm\varpi_{E_{2}}^{-1}, \pm 2\varpi_{E_{2}}^{-1}\} .
\]
Hence
\[
\ord\left(
X^{*}_{1}(\Lie(\alpha^{\vee})(1))
\right) = \ord(\varpi_{E_{2}}^{-1}) = -1/2 ,
\]
and $X_{1}^{*}$ satisfies {\bf (GE1)} in \cite[Definition~3.5.2]{Fintzen-IHES}.
Consider the element 
$\left(\begin{pmatrix}
	\varpi_{E_2} & 0 \\
	0 & 0 
\end{pmatrix}, 0\right) \in \Lie^*(G^1)_{x_0, -1/2}$, then
\[\ord\left( X_1^* \left(\begin{pmatrix}
	\varpi_{E_2} & 0 \\
	0 & 0 
\end{pmatrix}, 0\right) \right)=\ord(2)=0 ,
\]
thus  $X_1^*  \notin \Lie^*(G^1)_{x_0, -1/2+}$.
Moreover, $X_1^*  \in \Lie^*(G^1)_{x_0, -1/2}$, hence $X_{1}^{*}$ satisfies {\bf (GE0)} in \cite[Definition~3.5.2]{Fintzen-IHES}.

Since the only possible torsion prime for the dual root datum of $G^1$ and $G^2$ is 2, and since $p \neq 2$, by \cite[Lemma~8.1]{Yu} condition {\bf (GE2)} is also satisfied.
\end{proof}

As a consequence of the above lemma, the datum
\[
\Sigma = \bigl(
(G^0 \subset G^1 \subset G^2, M^0), (\tfrac14, \tfrac12, \tfrac12), (x_{0}, \{\iota\}), (K_{M^0}, \rho_{M^0}), (\phi_{0}, \phi_{1}, 1)
\bigr)
\]
satisfies the properties of \cite[(7.2)]{Kim-Yu}, in other words, it is a $G$-datum as in \cite[Definition~4.1.1]{HAIKY}.
Applying the construction of Kim and Yu in \cite[7.4]{Kim-Yu} to $\Sigma$, we obtain a compact, open subgroup $K$ of $G(F)$ and an irreducible smooth representation $\rho$ of $K$.

\begin{remark}
In \cite{HAIKY}, we twist the construction of Kim and Yu by a quadratic character $\epsilon_{x_{0}}^{\overrightarrow{G}}$ of $K^{0}/G^0(F)_{x_{0}, 0+} \simeq K_{M^0}/M^0(F)_{x_{0}, 0+}$ introduced in \cite{FKS}, see \cite[\S4.1]{HAIKY} for details.
In our case, we can compute $\epsilon_{x_{0}}^{\overrightarrow{G}}$ using \cite[Definition~3.1, Theorem~3.4]{FKS} as follows:
\[
\epsilon_{x_{0}}^{\overrightarrow{G}}\left(
(x, y, z) \bmod M^0(F)_{x_{0}, 0+}
\right) = \sgn_{\ff}(xy^{-1} \bmod \mathfrak{p}_{E_{2}}) = \sgn_{\ff}(xy \bmod \mathfrak{p}_{E_{2}})
\]
for $(x, y, z) \in K_{M^0}$, where $\sgn_{\ff}$ denotes the unique non-trivial quadratic character of $\ff^{\times} = \mathcal{O}^{\times}_{E_{2}}/(1+\mathfrak{p}_{E_{2}})$.
Since $-1 \in (\ff^{\times})^{2}$, the conditions in \eqref{descriptionofKM0} imply that 
\[
xy \bmod \mathfrak{p}_{E_{2}} = \pm (z \bmod \mathfrak{p}_{E_{4}})^{-2} \in (\ff^\times)^2.
\]
Thus, we obtain that $\epsilon_{x_{0}}^{\overrightarrow{G}}$ is trivial, and the twisted and non-twisted constructions agree in our case.
\end{remark}

According to
\cite[Theorem \ref{HAIKY-theoremstructureofheckeforKimYu}]{HAIKY},
we have an isomorphism of $\bC$-algebras
\[
\cH(G^0(F), (G^0(F)_{x_0}, \rho^0)) \isoarrow \cH(G(F), (K, \rho)).
\]
In the following section, we determine explicitly the structure of the Hecke algebras $\cH(G^0(F),(G^0(F)_{x_0}, \rho^0))$ and $\cH(G^0(F),(G^0(F)_{x_0,0}, \rho^0))$.

\section{Structure of the depth-zero Hecke algebra}
\label{sec:depth0}

In this section, we will study the Hecke algebra $\cH(G^0(F), (K^0, \rho^0))$ associated with the depth-zero type $(K^0, \rho^0)$.
We define the subgroup $N(\rho_{M^0})$
of the $F$-points of the normalizer $N_{G^0}(M^0)$ of $M^0$ in $G^0$ by
\[
N(\rho_{M^0}) :=
\left\{
n \in N_{G^0}(M^0)(F) \mid {^n\!K_{M^0}} = K_{M^0}, \, {^n\!\rho_{M^0}} = \rho_{M^0}
\right\}
\]
and write $W(\rho_{M^0}) := N(\rho_{M^0})/K_{M^0}$. 
We write 
\[
I_{G^0(F)}(\rho^0)
:=
\{
g \in G^0(F)
\mid
\Hom_{K^0 \cap ^g\! K^0}({^g\! \rho^0}, \rho^0) \neq \{0\}
\}.
\]
Then from 
\cite[Proposition~\ref{HAI-propproofofaxiombijectionofdoublecoset}
and
Corollary~\ref{HAI-corollarybijectiondoublecosetandweylgroup}]{HAI},
we have $N(\rho_{M^0}) \subset I_{G^0(F)}(\rho^0)$ and the inclusion induces a bijection
\[
W(\rho_{M^0}) \simeq K^0 \backslash I_{G^0(F)}(\rho^0) / K^0.
\]
In order to describe the groups $N(\rho_{M^0})$ and $W(\rho_{M^0})$ below, we define the element $\widetilde{s}$ of the group
\[
G^0(F) = \left(
\GL_{2}(E_{2}) \times \GL_{1}(E_{4})
\right) \cap \SL_{8}(F) \supset \SL_{2}(E_{2}) \times \SL_{1}(E_{4})
\]
by
$
\widetilde{s} = \left(
(\begin{smallmatrix}
\phm 0 & 1 \\
-1 & 0
\end{smallmatrix}), 1
\right) 
$.
Then we have $N_{G^0}(M^0)(F) = \{1, \widetilde{s}\} \ltimes M^0(F)$.
We note that the element $\widetilde{s}$ normalizes the groups $\widetilde{K}_{M^0}$ and $K_{M^0}$.
\begin{lemma}
\label{lemmasnormalizesthecharacterrhoM}
The element $\widetilde{s}$ normalizes the character $\rho_{M^0}$.
\end{lemma}
\begin{proof}
We have
\begin{align*}
^{\widetilde{s}}\!\widetilde\rho_{M^0} &= {^{\widetilde{s}}\!\Bigl(
1 \boxtimes (
\eta \circ N_{E_{2} / F}
) \boxtimes 1
\Bigr)} \\
&= 
(\eta \circ N_{E_{2} / F}) \boxtimes 1 \boxtimes 1 \\
&= \Bigl(
1 \boxtimes (\eta^{-1} \circ N_{E_{2} / F}) \boxtimes (\eta^{-1} \circ N_{E_{4} / F})
\Bigr) \otimes 
(\eta \circ \det),
\end{align*}
where $\det$ denotes the restriction of the determinant map $\GL_{8}(F) \to F^{\times}$ to the group $\widetilde{K}_{M^0}$.
Since the group $K_{M^0}$ is contained in the group $\SL_{8}(F)$, we have
\begin{equation}
\label{calculationofsrhoM}
^{\widetilde{s}}\!\rho_{M^0} = \Bigl(
1 \boxtimes (\eta^{-1} \circ N_{E_{2} / F}) \boxtimes (\eta^{-1} \circ N_{E_{4} / F})
\Bigr)\restriction_{K_{M^0}}.
\end{equation}
We will prove that
\begin{equation}
\label{etasquareE1trivial}
\eta^{2} \circ N_{E_{2} / F} \restriction_{\cO_{E_{2}}^{\times}} = 1
\end{equation}
and
\begin{equation}
\label{etaE2trivial}
\eta \circ N_{E_{4}/F} \restriction_{\cO_{E_{4}}^{\times}}  = 1.
\end{equation}
First, we will prove Equation~\eqref{etasquareE1trivial}.
The definition of $E_{2}$ implies that we have 
\[
N_{E_{2} / F}(\cO_{E_{2}}^{\times}) = (1 + \fp_{F}) \langle \zeta^{2} \rangle.
\]
Hence, Equation~\eqref{etasquareE1trivial} follows from the definition of $\eta$.
Similarly, we can prove Equation~\eqref{etaE2trivial} by using the definition of $\eta$ and the equation
\[
N_{E_{4} / F}(\cO_{E_{4}}^{\times}) = (1 + \fp_{F}) \langle \zeta^{4} \rangle.
\]

Combining equation~\eqref{calculationofsrhoM} with Equations~\eqref{etasquareE1trivial} and \eqref{etaE2trivial}, we obtain that
\begin{align*}
^{\widetilde{s}}\!\rho_{M^0} &= \Bigl(
1 \boxtimes (\eta^{-1} \circ N_{E_{2} / F}) \boxtimes (\eta^{-1} \circ N_{E_{4} / F})
\Bigr)\restriction_{K_{M^0}} \\
&= 
 \Bigl(
1 \boxtimes (\eta \circ N_{E_{2} / F}) \boxtimes 1
\Bigr)\restriction_{K_{M^0}} \\
&= \widetilde\rho_{M^0}\restriction_{K_{M^0}} = \rho_{M^0}.
\qedhere
\end{align*}
\end{proof}
\begin{proposition}
\label{propositionstructureofNrhoMSL8example}
We have
\[
N(\rho_{M^0}) = N_{G^0}(M^0)(F) = \{1, \widetilde{s}\} \ltimes M^0(F).
\]
\end{proposition}
\begin{proof}
The claim $N(\rho_{M^0}) \subset N_{G^0}(M^0)(F)$ follows from the definition of $N(\rho_{M^0})$.
We will prove the reverse inclusion.
According to Lemma~\ref{lemmasnormalizesthecharacterrhoM}, we have $\widetilde{s} \in N(\rho_{M^0})$.
Moreover, since $M^0$ is a torus, the conjugate action of $M^0(F)$ on $K_{M^0}$ is trivial.
Thus, we conclude that the group $M^0(F)$ normalizes the character $\rho_{M^0}$.
\end{proof}

To describe the structure of the group $W(\rho_{M^0})$,
we define the elements $\widetilde{s}' \in N(\rho_{M^0})$ and $\widetilde{z} \in M^0(F)$ by
$
\widetilde{s}' = \left(
\left(\begin{smallmatrix}
0 & \varpi_{E_{2}}^{-1} \\
-\varpi_{E_{2}} & 0
\end{smallmatrix}\right), 1
\right) 
$ 
and $\widetilde{z} = \left(
\zeta \varpi_{E_{2}}, \varpi_{E_{2}}, \varpi_{E_{4}}^{-2}
\right)$.
Let $s$, $s'$, and $z$ be the images of $\widetilde{s}$, $\widetilde{s}'$, and $\widetilde{z}$ in $W(\rho_{M^0})$, respectively.
When $K^0=G^0(F)_{x_0,0}$, we also set $\widetilde \epsilon_{M^0} := (-1, 1, 1) \in M^0(F)_{x_0} \smallsetminus M^0(F)_{x_0,0}$ and let $\epsilon_{M^0}$ denote the image of $\widetilde \epsilon_{M^0}$ in $W(\rho_{M^0})$.

\begin{corollary}
\label{cor:W-as-lattice-with-action}
We have
\(
W(\rho_{M^0}) = \{1, s\} \ltimes ( M^0(F) / K_{M^0} ).
\)
\end{corollary}

\begin{proof}
This is immediate from Proposition \ref{propositionstructureofNrhoMSL8example}.
\end{proof}
\begin{proposition}
\label{propositionstructureofWrhoMSL8example}
We have 
\[
W(\rho_{M^0}) =
\begin{cases}
 \langle
s, s'
\rangle \times \langle z \rangle \phantom{,\epsilon_{M^0}}
\simeq W_{\mathrm{aff}} (\widetilde A_1) \times \bZ & (K^0 = G^0(F)_{x_0}), \\
 \langle
s, s'
\rangle \times \langle z, \epsilon_{M^0} \rangle
\simeq W_{\mathrm{aff}} (\widetilde A_1) \times \bZ \times \bZ/2\bZ& (K^0 = G^0(F)_{x_0,0}),
\end{cases}
\]
where $W_{\mathrm{aff}}(\widetilde A_1)$ denotes the affine Weyl group of the affine root system of type $\widetilde A_1$, i.e., the affine Weyl group with two simple reflections and no relation between them.
\end{proposition}
\begin{proof}
First, we consider the case where $K^0 = G^0(F)_{x_0}$.
We define the isomorphism
\[
H_{M^0} \colon \tM^0(F) / \widetilde{K}_{M^0}  \rightarrow \mathbb{Z}^{3}
\]
by
\[
(x, y, z) \widetilde{K}_{M^0}  \mapsto (\ord^{\norm}_{E_{2}}(x), \ord^{\norm}_{E_{2}}(y), \ord^{\norm}_{E_{4}}(z)).
\]
The definition of $M^0$ implies that an element $(n_{1}, n_{2}, n_{3}) \in \mathbb{Z}$ is contained in the image of the subgroup
\[
M^0(F) / K_{M^0} \simeq M^0(F) \cdot \widetilde{K}_{M^0}  / \widetilde{K}_{M^0} 
\]
of $\tM^0(F) / \widetilde{K}_{M^0}$
if and only if
\begin{equation}
\label{conditionaboutnormmaps}
N_{E_{2}/F}(\varpi_{E_{2}}^{n_{1} + n_{2}}) \cdot N_{E_{4} / F}(\varpi_{E_{4}}^{n_{3}}) \in N_{E_{2}/F}(\cO_{E_{2}}^{\times}) \cdot N_{E_{4}/F}(\cO_{E_{4}}^{\times}).
\end{equation}
The definition of $E_{2}$ and $E_{4}$ implies that \eqref{conditionaboutnormmaps} is equivalent to the condition that
\[
\varpi_{F}^{n_{1} + n_{2} + n_{3}} \cdot \zeta^{n_{3}} \in (1 + \fp_{F}) \langle \zeta^{2} \rangle.
\]
Thus, we conclude that 
\begin{align*}
H_{M^0} \left(
M^0(F)/K_{M^0}
\right) &= \left\{
(n_{1}, n_{2}, n_{3}) \in \mathbb{Z} \mid n_{1} + n_{2} + n_{3} = 0 \:\text{and} \:  2 \mid n_{3}
\right\} \\
&= \langle 
(1, 1, -2), (1, -1, 0)
\rangle.
\end{align*}
Since
$
H_{M^0}(z) = (1, 1, -2)
$,
$H_{M^0}(s s') = (1, -1, 0)$, and $H_{M^0}$ is an isomorphism, we have that 
\[
M^0(F)/K_{M^0} = \langle 
ss', z
\rangle. 
\]
From Corollary \ref{cor:W-as-lattice-with-action},
we thus obtain that
\[
W(\rho_{M^0}) = \{1, s\} \ltimes \left(
M^0(F) / K_{M^0}
\right) = \langle
s, s', z
\rangle.
\]
One can check that the subgroup $\langle
s, s'
\rangle$ of $W(\rho_{M^0})$ is isomorphic to the affine Weyl group of type $\widetilde{A_{1}}$, and that the element $z$ has infinite order, commutes with the elements $s$ and $s'$, and no non-trivial power of $z$ is contained in the span of $s$ and $s'$.

Next, we consider the case where $K^0 = G^0(F)_{x_0,0}$.
In this case, we have
\[
M^0(F)/K_{M^0} = M^0(F)/M^0(F)_{x_0,0} = \left(
M^0(F)/M^0(F)_{x_0}
\right) \times \langle \epsilon_{M^0} \rangle.
\]
Noting that $\epsilon_{M^0}$ commutes with $s$ and $s'$, the claim follows from the first case.
\end{proof}

\begin{lemma}
\label{lemmacanttakenznscommute}
Let $n_{s}, n_{z} \in N(\rho_{M^0})$
denote lifts of $s$ and $z$.
Then, we have $[n_{s}, n_{z}] \in K_{M^0} \smallsetminus \ker \rho_{M^0}$.
In particular, we have $n_{s} n_{z} \neq n_{z} n_{s}$.
\end{lemma}
\begin{proof}
We write $n_{s} = \widetilde{s} k$ and $n_{z} = \widetilde{z} k'$
for some $k, k' \in K_{M^0}$.
Then, we have
\begin{align*}
[n_{s}, n_{z}] &= [\widetilde{s} k, \widetilde{z} k'] \\
&= [\widetilde{s} k, \widetilde{z}] \cdot ^{\widetilde{z}}\![\widetilde{s} k, k'] \\
&= {^{\widetilde{s}}[k, \widetilde{z}]} \cdot [\widetilde{s}, \widetilde{z}] \cdot ^{\widetilde{z}}\![\widetilde{s} k, k'].
\end{align*}
Since $\widetilde{s}, \widetilde{z} \in N(\rho_{M^0})$ normalize $\rho_{M^0}$, we have
\[
^{\widetilde{s}}[k, \widetilde{z}], ^{\widetilde{z}}\![\widetilde{s} k, k'] \in \ker \rho_{M^0}.
\]
On the other hand, we have 
\[
[\widetilde{s}, \widetilde{z}] = (\zeta^{-1}, \zeta, 1) \in K_{M^0} \smallsetminus \ker \rho_{M^0}
\]
since
\[
\rho_{M^0}((\zeta^{-1}, \zeta, 1)) = \left(
\eta \circ N_{E_{2}/F}
\right)(\zeta) = \eta(\zeta^{2}) = -1.
\]
Thus, we conclude that $[n_{s}, n_{z}] \in K_{M^0} \smallsetminus \ker \rho_{M^0}$.
\end{proof}

\begin{corollary}
\label{corollaryrhoMdoesnotextend}
The character $\rho_{M^0}$ does not extend to the group $N(\rho_{M^0})$.
\end{corollary}
\begin{proof}
Suppose that $\rho_{M^0}$ extends to a character $\rho_{M^0}^{\dagger}$of $N(\rho_{M^0})$.
Then, since $\widetilde{s}, \widetilde{z} \in N(\rho_{M^0})$,
we have $[\widetilde{s}, \widetilde{z}] \in \ker \rho_{M^0}^{\dagger}$,
which contradicts Lemma~\ref{lemmacanttakenznscommute}.
\end{proof}

Our decomposition of $W(\rho_{M^0})$ given in
Proposition \ref{propositionstructureofWrhoMSL8example}
gives rise to a length function on this group, the standard length function on extended affine Weyl groups. More precisely, we start with the length function $\plength$ on $\langle
s, s'
\rangle$ with respect to the generators $\{s, s'\}$ of $\langle
s, s'
\rangle$.
We extend $\plength$ to $W(\rho_{M^0})$ by
\[
\begin{cases}
\plength(w z^{n}) := \plength(w) & \text{ when } K^0 = G^0(F)_{x_0}, \\
\plength(w z^{n} \epsilon_{M^0}^{t}) := \plength(w) & \text{ when } K^0 = G^0(F)_{x_0,0}
\end{cases}
\]
for $w \in \langle
s, s'
\rangle$, $n \in \mathbb{Z}$, and $t \in \{0, 1\}$.

\begin{remark}
\label{rmk:no-good-coset-reps}
Suppose that $K^0 = G^0(F)_{x_0,0}$.
According to \cite[Proposition~5.2]{Morris}, we can take a lift $n_{w} \in N(\rho_{M^0})$ for each $w \in W(\rho_{M^0})$ such that if $\plength(w_{1} w_{2}) = \plength(w_{1}) + \plength(w_{2})$, then we have $n_{w_{1} w_{2}} = n_{w_{1}} n_{w_{2}}$.
However, this statement is false in general,
and Lemma~\ref{lemmacanttakenznscommute} provides a counterexample. 
The failure of \cite[Proposition~5.2]{Morris} does not affect Morris's main result \cite[Theorem 7.12]{Morris} as his proof can easily be adapted to circumvent the use of such good coset representatives. Alternatively, the recent proof of the more general result \cite[Section \ref{HAI-sec:depth-zero}]{HAI} also does not rely on a choice of representatives.
On the other hand, \cite[Remark 7.12(a)]{Morris}, which states that the the $2$-cocycle $\muT$ is trivial if the representation $\rho^{0}$ is a character, does depend on such representatives, and the example covered in this paper shows that \cite[Remark 7.12(a)]{Morris} is not true in general (see Corollary~\ref{corollaryanexampleofheckealgebrawithnon-trivialtwococycle} below).
\end{remark}

Although 
Proposition \ref{propositionstructureofWrhoMSL8example}
decomposes $W(\rho_{M^0})$ into a product
of an affine Weyl group $\langle s,s' \rangle$
and a complement ($\langle z \rangle$ and $\langle z, \epsilon_{M^0} \rangle$ for $K^0 = G^0(F)_{x_0}$ and $K^0 = G^0(F)_{x_0,0}$, respectively)
the decomposition of $W(\rho_{M^0})$ provided in 
\cite{HAI} (and also in \cite[7.3]{Morris}) is different, and comes from a different length function, where more elements have length zero,
which is denoted by $\flength$ and defined in \cite[Definition \ref{HAI-definitionlength}]{HAI}.
The subgroup $\Wzeroz$ of $W(\rho_{M^0})$ is defined by
\[
	\Wzeroz := \Bigl\{
	w \in W(\rho_{M^0}) \Bigr. \,\Bigl | \, \flength(w) = 0
	\Bigr\}.
\]

\begin{proposition}
\label{propositionWheartisOmegainSL8example}
We have $W(\rho_{M^0}) = \Omega(\rho_{M^0})$.
\end{proposition}

That is, all elements of $W(\rho_{M^0})$ have length zero
with respect to $\flength$.

\begin{proof}
For each $w \in W(\rho_{M^0})$, we fix a non-zero element $\varphi_{w} \in \cH(G^0(F), (K^0, \rho^0))$
with support in ${K}^0 w {K}^0$.
To prove the proposition, it suffices to show that for any $w_{1}, w_{2} \in W(\rho_{M^0})$, we have $\varphi_{w_{1}} * \varphi_{w_{2}} \in \mathbb{C} \cdot \varphi_{w_{1} w_{2}}$.
According to Proposition~\ref{propositionstructureofWrhoMSL8example}, we have
$
W(\rho_{M^0}) = \langle
s, s'
\rangle \times \langle z \rangle
$ or $
W(\rho_{M^0}) = \langle
s, s'
\rangle \times \langle z, \epsilon_{M^0} \rangle
$,
and we can check easily that if $w_{1}, w_{2} \in W(\rho_{M^0})$ satisfy $\plength(w_{1} w_{2}) = \plength(w_{1}) + \plength(w_{2})$, then we have 
\[
\widetilde{K}^0 w_{1} \widetilde{K}^0 w_{2} \widetilde{K}^0 = \widetilde{K}^0 w_{1} w_{2} \widetilde{K}^0.
\]
Since we have $\widetilde{K}^0 = \widetilde{K}_{M^0} \cdot K^0$, and the group $N(\rho_{M^0})$ normalizes the group $\widetilde{K}_{M^0}$, we also obtain that 
\begin{align*}
K^0 w_{1} K^0 w_{2} K^0 & \subseteq \widetilde{K}^0 w_{1} \widetilde{K}^0 w_{2} \widetilde{K}^0 \cap \SL_{8}(F) \\
&= \widetilde{K}^0 w_{1} w_{2} \widetilde{K}^0 \cap \SL_{8}(F) \\
&= \widetilde{K}^0 w_{1} w_{2} \widetilde{K}_{M^0} \cdot K^0 \cap \SL_{8}(F) \\
&= \widetilde{K}^0 \cdot  \widetilde{K}_{M^0}  w_{1} w_{2} K^0 \cap \SL_{8}(F) \\
&= \widetilde{K}^0 w_{1} w_{2} K^0 \cap \SL_{8}(F) \\
&= \left(
\widetilde{K}^0 \cap \SL_{8}(F)
\right) w_{1} w_{2} K^0 \\
&= K^0 w_{1} w_{2} K^0.
\end{align*}
In particular, in this case, we obtain that $\varphi_{w_{1}} * \varphi_{w_{2}} \in \mathbb{C} \cdot \varphi_{w_{1} w_{2}}$.
Thus, to prove the proposition, it now suffices to show that $\varphi_{s} * \varphi_{s} \in \mathbb{C} \cdot \varphi_{1}$ and $\varphi_{s'} * \varphi_{s'} \in \mathbb{C} \cdot \varphi_{1}$.
Similar calculations as above imply that $K^0 s K^0 s K^0 = K^0 \cup K^0 s K^0$ and $K^0 s' K^0 s' K^0 = K^0 \cup K^0 s' K^0$.
Hence, we obtain that
\[
\varphi_{s} * \varphi_{s} \in \mathbb{C} \cdot \varphi_{s} \oplus \mathbb{C} \cdot \varphi_{1}
\qquad
\text{and}
\qquad
\varphi_{s'} * \varphi_{s'} \in \mathbb{C} \cdot \varphi_{s'} \oplus \mathbb{C} \cdot \varphi_{1}.
\]
Thus, it suffices to prove that $\left(
\varphi_{s} * \varphi_{s}
\right)(\widetilde{s}) = \left(
\varphi_{s'} * \varphi_{s'}
\right)(\widetilde{s}') = 0$.
We take a set of representatives for 
\(
K^0 / \left(
K^0 \cap {^{\widetilde{s}}\!K^0}
\right)
\)
as
\(
\left\{
u(x) \mid x \in \cO_{E_{2}} / \fp_{E_{2}}
\right\}
\),
where
$
u(x) = \left(
\begin{pmatrix}
1 & x \\
0 & 1
\end{pmatrix}, 1
\right)
$.
Then, we can calculate the convolution product $\left(
\varphi_{s} * \varphi_{s}
\right)(\widetilde{s})$ as
\begin{align*}
\left(
\varphi_{s} * \varphi_{s}
\right)(\widetilde{s}) 
&= \sum_{h \in K^0 \widetilde{s} K^0 / K^0} \varphi_{s}(h) \cdot \varphi_{s}(h^{-1}\widetilde{s}) \\
&= \sum_{k \in K^0 / \left(
K^0 \cap {^{\widetilde{s}}\!K^0}
\right)
} \varphi_{s}(k\widetilde{s}) \cdot \varphi_{s}(\widetilde{s}^{-1} k^{-1} s) \\
&= \sum_{x \in \cO_{E_{2}} / \fp_{E_{2}}}  \varphi_{s}(u(x) \widetilde{s}) \cdot \varphi_{s}(\widetilde{s}^{-1} u(-x) \widetilde{s}).
\end{align*}
For $x \in \cO_{E_{2}}$, we have
\begin{align*}
\widetilde{s}^{-1} u(-x) \widetilde{s} 
=
\left(
\begin{pmatrix}
1 & 0 \\
x & 1
\end{pmatrix}, 1
\right).
\end{align*}
Hence, $\widetilde{s}^{-1} u(-0) \widetilde{s} \not \in K^0 \widetilde{s} K^0$ and for $x \in \cO_{E_{2}}^{\times}$, we have
\begin{align*}
\widetilde{s}^{-1} u(-x) \widetilde{s} &= 
\left(
\begin{pmatrix}
-x^{-1} & -1 \\
0 & -x 
\end{pmatrix}
\cdot
\widetilde{s}
\cdot
\begin{pmatrix}
1 & x^{-1} \\
0 & 1
\end{pmatrix}, 1
\right).
\end{align*}
Hence, the definition of $\rho^0$ implies that
\begin{align*}
\left(
\varphi_{s} * \varphi_{s}
\right)(\widetilde{s}) &= \sum_{x \in \cO_{E_{2}} / \fp_{E_{2}}}  \varphi_{s}(u(x) \widetilde{s}) \cdot \varphi_{s}(\widetilde{s}^{-1} u(-x) \widetilde{s}) \\
&= \varphi_{s}(\widetilde{s})^{2} \sum_{x \in \cO_{E_{2}}^{\times} /\left(1+ \fp_{E_{2}}\right)} \left(
\eta \circ N_{E_{2}/F}
\right)(-x) \\
&= \varphi_{s}(\widetilde{s})^{2} \sum_{x \in \cO_{F}^{\times} /\left(1+ \fp_{F}\right)} \eta(x^{2}) = 0,
\end{align*}
where the last equality follows from the fact that the restriction of the character $\eta^{2}$ to $\cO_{F}^{\times}$ is non-trivial.
Similarly, we can prove that $\left(
\varphi_{s'} * \varphi_{s'}
\right)(\widetilde{s}') = 0$.
\end{proof}

We fix a family
$
\cT =
\left\{
T_{n} \in \Hom_{K_{M^0}}\left(
^n\!\rho_{M^0}, \rho_{M^0}
\right)
\right\}_{n \in N(\rho_{M^0})}
$
as in \cite[Choice \ref{HAI-choice:tw}]{HAI}
and define the $2$-cocycle 
\[
\muT \colon W(\rho_{M^0}) \times W(\rho_{M^0}) \rightarrow \bC^{\times}
\]
as in \cite[Notation \ref{HAI-notationofthetwococycle}]{HAI}.

\begin{corollary}
\label{corollaryanexampleofheckealgebrawithnon-trivialtwococycle}
We have an isomorphism
\[
\cH(G^0(F), (K^0, \rho^0)) \simeq \mathbb{C}[W(\rho_{M^0}), \muT],
\]
and the $2$-cocycle $\muT$ is non-trivial.
\end{corollary}
\begin{proof}
The corollary follows from
\cite[Theorem \ref{HAI-thm:isomorphismtodepthzero}]{HAI},
Corollary~\ref{corollaryrhoMdoesnotextend},
and
Proposition~\ref{propositionWheartisOmegainSL8example}.
\end{proof}

\bibliographystyle{amsalpha}
\bibliography{ourbib}
\end{document}